OpenStudy (kimberlyevens):

Will MEDAL and FAN!! 4. Is triangle ABC with vertices A(-1,4), B(3,1), and C(0,-3) a right triangle? Explain your answer using the slopes of the segments that form the sides of the triangle.

OpenStudy (campbell_st):

well plot the points then find the slope of the line segments lastly multiply pairs of slopes the pair that multiply to -1 are perpendicular... and hence right angled

OpenStudy (kimberlyevens):

@waterineyes

OpenStudy (anonymous):

Just follow the steps my friend campbell has given you above..

OpenStudy (kimberlyevens):

where do i plot the points

OpenStudy (anonymous):

On a graph..

OpenStudy (anonymous):

Here, just a rough sketch will work: 

OpenStudy (kimberlyevens):

thanks

OpenStudy (kimberlyevens):

i ploted the points how do i find the slope of the lie segment

jimthompson5910 (jim_thompson5910):

you use the slope formula m = (y2 - y1)/(x2 - x1) to find the slope of the line through any two points

jimthompson5910 (jim_thompson5910):

if you can show two slopes m1 and m2 multiply to -1, like so m1*m2 = -1 then you have shown you have a right triangle

OpenStudy (kimberlyevens):

this is what i got when i ploted the points @jim_thompson5910

jimthompson5910 (jim_thompson5910):

are you able to find the slopes of lines: AC, BC, AB ?

OpenStudy (kimberlyevens):

no thats what i need help with

jimthompson5910 (jim_thompson5910):

let's find the slope of AB A(-1,4), B(3,1)

jimthompson5910 (jim_thompson5910):

plug the coordinates into m = (y2 - y1)/(x2 - x1) in this case x1 = -1 y1 = 4 x2 = 3 y2 = 1

jimthompson5910 (jim_thompson5910):

what slope value do you get?

OpenStudy (kimberlyevens):

give me a minute i have take my dog outside

jimthompson5910 (jim_thompson5910):

ok

OpenStudy (kimberlyevens):

so this is what i got m=(4*2-1*1)/(-1*2-3*1) m=-1.75 is that right? @jim_thompson5910

jimthompson5910 (jim_thompson5910):

the y2 does NOT mean y times 2 y2 is simply a label to tell the difference from y1 both are y values (just different y values)

jimthompson5910 (jim_thompson5910):

y1 is the first y value y2 is the second y value y2 - y1 tells us to subtract the two

jimthompson5910 (jim_thompson5910):

same goes for x1 and x2

OpenStudy (kimberlyevens):

ok give me a minute